OpenStudy (anonymous):

WILL FAN AND MEDAL!!!! Three years ago, Jolene bought $750 worth of stock in a software company. Since then th value of her purchase has been increasing at an average of 12 3\5% per year. How much is the stock worth now. 3\5 is a fraction

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@mathmale

OpenStudy (retireed):

F = P e^rt r = (12 3/5) / 100 = 12.6/100 = 0.126 = 750 e^(0.126*3) = _______ ??